OpenStudy (anonymous):

how to find moment generating function? Anyone can explain with the example that if X denotes the out come when fair die is tossed find MGF, mean and variance.

OpenStudy (anonymous):

@hartnn buddy can you help me with this?

OpenStudy (anonymous):

mathworld.wolfram.com/Moment-GeneratingFunction.html

OpenStudy (anonymous):

@Hero can you help me with this?

OpenStudy (anonymous):

@zepdrix can you help me?

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

anyone? O_________O

OpenStudy (amistre64):

https://onlinecourses.science.psu.edu/stat414/node/52 im reading thru this to see if i can learn about MGFs

OpenStudy (amistre64):

\[\mu=E(X)=\frac16\]might be needed

OpenStudy (amistre64):

\[E(X^2)=\sigma^2+\mu^2\]seems line another

OpenStudy (anonymous):

hey @amistre64 check this from same site https://onlinecourses.science.psu.edu/stat414/node/72

OpenStudy (amistre64):

already there :)

OpenStudy (amistre64):

p = 1/6, 1-p = 5/6

OpenStudy (anonymous):

M(t)=E(etX)=∑ e^tx * f(x) I am not getting whose summation is to be taken?

OpenStudy (amistre64):

im not sure at the moment why the e^(tx) is important

OpenStudy (anonymous):

o_o because it is in the formula for mgf= M(t)

OpenStudy (anonymous):

still if you have something else please go on..

OpenStudy (amistre64):

yeah, i see that :) i just like to know why its there ... the function is the probability distribution; which in this case is uniform. f(x) = 1/6, from 0 to 6  and zero everywhere else

OpenStudy (anonymous):

yep thats correct

OpenStudy (amistre64):

or could we go with the binomial of p=1/6, 1-p = 5/6 ?

OpenStudy (anonymous):

anything, however you want. I just want to learn you get the MGF here I'll understand the thing.

OpenStudy (amistre64):

well, from what i see, we take that e^(tx) and multiply it to our probability distribution function. and simplify

OpenStudy (amistre64):

\[f(x)=\frac16~:~(0,6)\] but thats a continuous random variable setup .... seeing how there is no real way to get x=3.5, this might not be the best setup for f(x)

OpenStudy (anonymous):

:( what to do then?

OpenStudy (amistre64):

hmmm, if we wanted to know the probability of getting exactly one 3, out of 5 tosses of the die, we would run a binomial setup ffffs fffsf ffsff fsfff sffff these are the number of ways to get exactly 1 "value" out of 5 tosses; s = success, f = fail 5 * (s)^1* (f)^4\[\binom{5}{1}(\frac16)^1(\frac56)^4\]

OpenStudy (amistre64):

so it depends on what it is we are trying to determine here

OpenStudy (amistre64):

\[\sum_{x=0}^6~ \frac16e^{tx}\]doesnt seem to mean much to me

OpenStudy (anonymous):

yup this one I know... and I think we are going correct as per the video M(t) = ( (1-p) + p* e^t ) ^n so here M(t) = ( 5/6 + (1/6)* e^t )^n must be the MGF

OpenStudy (amistre64):

i agree

OpenStudy (anonymous):

then I think 1st derivative of mgf with t=0 give mean

OpenStudy (anonymous):

dont know what to do. am I right???

OpenStudy (amistre64):

derivatives with respect to "t", yes ... evaluated at t=0

OpenStudy (anonymous):

what do you get?? I got 1 as mean o.O. is is right?

OpenStudy (amistre64):

assuming we want n=1\[1[\frac56+\frac16e^{0}]^{0}~\frac16e^0\] \[1(1)\frac16(1)=\frac16\]

OpenStudy (anonymous):

ahan thats correct general formula for mean is n*p here n is 1 and p is 1/6 so mean will be 1/6. Thats absolutely correct. :)

OpenStudy (anonymous):

now can you find variance in same way ?

OpenStudy (amistre64):

variance is the second derivative w.r.t.t \[M''(t)=n(n-1)[(1-p)+pe^{t}]^{n-2}~p^2e^{2t}+n[(1-p)+pe^{t}]^{n-1}~pe^{t}\] \[M''(0)=n(n-1)[(1-p)+p]^{n-2}~p^2+n[(1-p)+p]^{n-1}~p\]

OpenStudy (anonymous):

I think variance should be 2nd derivative minus mean square so in the above equation as we have n=1 1st term will be zero and 2nd term 1/6 and variance will be 1/6- (1/6)^2 = 5/36

OpenStudy (anonymous):

correct me if I am wrong

OpenStudy (amistre64):

variance = np(1-p) = 5/36 .. yes, so i have to consider it i did the derivative correctly

OpenStudy (anonymous):

:D yup thats correct. we have done it. Thanks buddy.

OpenStudy (amistre64):

youre welcome